1. Air flow steadily at the rate of 0.95 kg/s through air compressor, entering at 7.5 m/s speed, 95 Kpaa

pressure and 0.95 m 3 /kg specific volume and leaving 20 m/s, 725 kpaa and 0.19 m 3 /kg. the internal

energy of the air leaving is 100 KJ/kg greater than that of the air entering. Cooling water in the

compressor jacket absorbs heat from the air at the rate of 60 kJ/s. compute the steady flow work in kW.


1
Expert's answer
2021-09-15T16:16:47-0400

"m= 0.95 \;kg/sec \\
v_1 = 0.95 \;m^3/kg \\
v_2= 0.19 \;m^3/kg \\
V_1 = 7.5 \;m/s \\
V_2 = 20 \;m/s \\
p_1 = 95 \;kPa = 95 \times 10^3 \;Pa \\
p_2 = 725 \;kPa = 725 \times 10^3 \;Pa \\
Q= -60 \;kJ/s \\
Δu=100 \;kJ/kg \\
wf_1 = p_1v_1 \\
= 95 \;kPa \times \frac{0.95 \; m^3}{kg} = 90.25 \;kJ/kg \\
wf_2 = p_2v_2 \\
= 725 \;kPa \times \frac{0.19 \; m^3}{kg} = 137.75 \;kJ/kg \\
KE_1 = \frac{V_1^2}{2k} \\
= \frac{(7.5 \; m/s)^2}{2 \times \frac{1 kg \;m \cdot \;m}{N \cdot s^2}} \\
= 0.028125 \;kJ/kg \\
KE_2 = \frac{V^2_2}{2k} \\
= \frac{(20 \;m/s)^2}{2 \times \frac{1 kg \;m \cdot \;m}{N \cdot s^2}} \\
= 0.2 \; kJ/kg \\
E_{in} = E_{out} \\
u_1 + wf_1 +KE_1+Q = w +u_2 +wf_2 +KE_2 \\
w = Δu +wf_2 +KE_2 - wf_1 -KE_1 -Q \\
= 100 \;kJ/kg + 137.75 \;kJ/kg+ 0.2 \; kJ/kg -90.25 \;kJ/kg -0.028 \;kJ/kg -(-60 \;kJ/s) \\
= 147.672 \;kJ/kg \times 0.95 \;kg/s + 60 \;kJ/s \; \\
= 140.288 \;kJ/s + 60 \;kJ/s \\
= 200.28 \;kJ/s \\
= 200.28 \;kW"